OpenStudy (anonymous):

What is the first step to solve the system using elimination: 4x+y=6 6x+y=2

OpenStudy (anonymous):

4X+Y=6 6X+Y=2 BY MULTIPLYING (-) SIGN IN THE SECOND EQUATION WE GET, 4X+Y=6 -6X-Y=-2 NOW THE Y TERM IS CANCELLED..... SO THE EQUTION OBTAINED IS...... -2X=4 SO, X=-2